OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of ?

OpenStudy (noelgreco):

The question is incomplete. You either need the slope or another point.

OpenStudy (anonymous):

The slope is 3/5

OpenStudy (anonymous):

y - y1 = m*(x-x1) m is slope

OpenStudy (anonymous):

^ I know this

OpenStudy (anonymous):

plug in the values then

OpenStudy (anonymous):

I did. 5x – 3y = 19 5x – 3y = –19 3x – 5y = 19 3x – 5y = –19 The answer choices aren't what I am getting!

OpenStudy (anonymous):

show me how u solved it

OpenStudy (anonymous):

y-y1=m(x-x1) plug in the values y+2=3/5(x-3) don't know where to go from there

OpenStudy (anonymous):

y+2=3/5(x-3) 5(y+2)=3(x-3) 5y + 10 = 3x - 9 10+ 9 = 3x - 5y 3x - 5y = 19 C

OpenStudy (anonymous):

Thanks!